### v1

@@ 4 Stationary Points and Irrelevant Feature Elimination
$\begin{split}\mathrm{D}\mathcal{J}(\beta,\lambda)[v]&=-\lim_{s\to 0^{+}}\frac{1}{s}(\mathcal{J}(\beta^{(s)},\lambda)-\mathcal{J}(\beta,\lambda))\\ &\leq-\lambda\liminf_{s\to 0^{+}}\frac{1}{s}(\left\|{P_{s}f_{\beta^{(s)},\lambda}}\right\|_{H}^{2}-\left\|{f_{\beta^{(s)},\lambda}}\right\|_{H}^{2})\\ &\leq-2\pi^{2}\lambda\liminf_{s\to 0^{+}}\int\frac{|\widehat{f}_{\beta^{(s)},\lambda}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}d\omega.\end{split}$
Since $\beta^{(s)}\to\beta$ by Lemma 9, we get $\left\|{f_{\beta^{(s)},\lambda}-f_{\beta,\lambda}}\right\|_{H}\to 0$, as $s\to 0^{+}$. Therefore, for every $R>0$, we apply Lebesgue’s dominated convergence theorem to get
$\begin{split}\liminf_{s\to 0^{+}}\int\frac{|\widehat{f}_{\beta^{(s)},\lambda}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}d\omega&\geq\liminf_{s\to 0^{+}}\int_{\left\|{\omega}\right\|\leq R}\frac{|\widehat{f}_{\beta^{(s)},\lambda}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}d\omega\\ &=\int_{\left\|{\omega}\right\|\leq R}\frac{|\widehat{f}_{\beta,\lambda}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}d\omega.\end{split}$
Since the result holds for every $R>0$, we deduce
$\mathrm{D}\mathcal{J}(\beta,\lambda)[v]\leq-2\pi^{2}\lambda\int\frac{|\widehat{f}_{\beta,\lambda}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}d\omega.$
The right-hand side equals zero only in one of two cases: either $f_{\beta,\lambda}$ is identically zero, which implies $\mathcal{J}(\beta,\lambda)=\mathbb{E}[Y^{2}]$, or $\beta_{S_{*}^{c}}=0$, meaning $\beta$ has no support outside the relevant coordinates. ∎
@@ 5 Stationary Points and Relevant Feature Recovery
The answer to $(\mathsf{Q4})$ is subtle and depends not only on the data distribution but, more crucially, on the choice of kernel $K$. The kernel defines the RKHS $H$ and shapes the nonconvex landscape of $\mathcal{J}$, thereby influencing which features can be identified at stationary points.
Our analysis shows that $\ell_{1}$-type RKHS (Example 4) are particularly effective for this purpose, enabling recovery of relevant features under milder distributional assumptions than commonly used radial kernel RKHS (Example 3) such as the Gaussian RKHS. We outline the underlying intuition in Section 5.1, and briefly recall the kernel class here.
@@ 5.1 Why the Laplace Kernel? A First Indication
We aim to understand when stationary points of $\mathcal{J}$ recover the relevant features—this is the central question posed in $(\mathsf{Q4})$. As we will see, the answer depends not only on the data distribution but also on the choice of kernel $K$, which fundamentally shapes the optimization landscape.
To build intuition for this kernel dependence, we consider a simpler diagnostic question:
@@
This question $(\mathsf{Q5})$ serves as a basic test of whether the kernel supports signal detection. Question (i) is the most basic: it asks whether the trivial solution $\beta=0$ can be ruled out a stationary point. Question (ii) is stronger—it asks whether all stationary points capture some relevant feature. The two questions in $(\mathsf{Q5})$ are in fact equivalent under the additional assumptions of Theorem 7, since Theorem 7 ensures that every stationary point satisfies $\mathop{\rm supp}(\beta)\subseteq S_{*}$.
@@ 5.1.1 Comparison to Radial Basis Kernels
$\mathcal{J}(se_{i},\lambda)-\mathcal{J}(0,\lambda)=\int_{0}^{s}\mathrm{D}\mathcal{J}(te_{i},\lambda)[e_{i}]dt=\int_{0}^{s}tG_{i}(t)dt=\frac{s^{2}}{2}G_{i}(0)+o(s^{2})~~(s\to 0).$
@@ 5.2.3 Error Control
$\begin{split}&\int_{0}^{\infty}\int_{\mathbb{R}^{d}}|\mathbb{E}[R_{\beta,\omega,\lambda}(X,Y)]|^{2}q_{t}(\omega)\,d\omega\,\mu(dt)\\ &=\mathbb{E}\left[r_{\beta,\lambda}(X,Y)r_{\beta,\lambda}(X^{\prime},Y^{\prime})\cdot\int_{0}^{\infty}\int_{\mathbb{R}^{d}}e^{-2\pi i\langle\omega,\beta\circ(X-X^{\prime})\rangle}q_{t}(\omega)\,d\omega\,\mu(dt)\right]\\ &=\mathbb{E}\left[r_{\beta,\lambda}(X,Y)r_{\beta,\lambda}(X^{\prime},Y^{\prime})\cdot K_{\beta}(X,X^{\prime})\right].\end{split}$
@@
$\lambda\cdot|\mathrm{D}\mathcal{J}(\beta,\lambda)[e_{i}]-\mathcal{S}_{i}(\beta,\lambda)|=\Delta_{1}+\Delta_{2}+\Delta_{3}$
@@
$\begin{split}\Delta_{1}&=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\mathbb{E}[R_{\beta,\omega,\lambda}(X,Y)]\mathbb{E}[\overline{R_{\beta,\omega,\lambda}(X^{\prime},Y^{\prime})}|X_{i}-X_{i}^{\prime}|]~tq_{t}(\omega)d\omega\mu(dt)\\ \Delta_{2}&=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\mathbb{E}[\overline{R_{\beta,\omega,\lambda}(X,Y)}]\mathbb{E}[R_{\beta,\omega,\lambda}(X^{\prime},Y^{\prime})|X_{i}-X_{i}^{\prime}|]~tq_{t}(\omega)d\omega\mu(dt)\\ \Delta_{3}&=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\mathbb{E}[R_{\beta,\omega,\lambda}(X,Y)]\mathbb{E}[\overline{R_{\beta,\omega,\lambda}(X^{\prime},Y^{\prime})}]~tq_{t}(\omega)d\omega\mu(dt)\end{split}$
The measure $tq_{t}(\omega)d\omega\mu(dt)$ is finite with $\int tq_{t}(\omega)d\omega\mu(dt)\leq M_{\mu}$. Applying the Cauchy–Schwarz inequality and invoking Lemmas 18 and 19, we obtain that the error terms $|\Delta_{1}|,|\Delta_{2}|$ are bounded by $C\sqrt{\lambda}$ and $|\Delta_{3}|$ is bounded by $C\lambda$, where $C<\infty$ depends only on $M_{X},M_{Y},M_{\mu}$.
@@ 5.3.1 Decomposition of Directional Derivative under Functional ANOVA
In particular, the right-hand side—and hence the covariance term—is independent of $\lambda$.
@@ 6 Open Problems
  • (Dependence) In the context of predictive feature recovery, our results assume independence between covariates. As explained earlier, this independence is mainly used for attributing prediction signals to individual features and feature subsets. Extending the analysis to dependent designs is particularly interesting.
@@ A.2.2 The functional $\mathfrak{L}$ and its Differentiability
$\mathrm{D}\mathfrak{L}(g,\beta,\lambda)[v]:=\lim_{s\to 0^{+}}\frac{1}{s}(\mathfrak{L}(g,\beta+sd,\lambda)-\mathfrak{L}(g,\beta,\lambda)).$ (A.3)

### v2

@@ 4 Stationary Points and Irrelevant Feature Elimination
$\begin{split}\mathrm{D}\mathcal{J}(\beta,\lambda)[v]&=\lim_{s\to 0^{+}}\frac{2}{s}(\mathcal{J}(\beta^{(s)},\lambda)-\mathcal{J}(\beta,\lambda)))\\ &\leq\lambda\cdot\limsup_{s\to 0^{+}}\frac{2}{s}(\left\|{P_{s}f_{\beta,\lambda}}\right\|_{H}^{2}-\left\|{f_{\beta,\lambda}}\right\|_{H}^{2})\\ &\leq-4\pi^{2}\lambda\cdot\int\frac{|\widehat{f}_{\beta,\lambda}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}d\omega.\end{split}$
The right-hand-side vanishes if and only if
$\frac{|\widehat{f}_{\beta,\lambda}|(\omega)^{2}}{k(\omega)}\cdot\|(\beta\circ\omega)_{S_{*}^{c}}\|^{2}_{\Sigma}=0~~~\text{a.e. $\omega$}.$
Since $\Sigma={\rm Cov}(X_{S_{*}^{c}})$ is positive definite, this happens if and only if: (i) either $\beta_{S_{*}^{c}}=0$, so that $\Pi_{S_{*}^{c}}\beta=0$; or (ii) $\widehat{f}_{\beta,\lambda}(\omega)=0$ for almost every $\omega$, in which case $f_{\beta,\lambda}\equiv 0$ and hence $\mathcal{J}(\beta,\lambda)=\mathbb{E}[Y^{2}]$. ∎
@@ 5 Stationary Points and Relevant Feature Recovery
The answer to $(\mathsf{Q4})$ is subtle and depends not only on the data distribution but, more crucially, on the choice of kernel $K$. Our analysis shows that $\ell_{1}$-type RKHS (Example 4) are particularly effective for this purpose, enabling recovery of relevant features under milder distributional assumptions than commonly used radial kernel RKHS (Example 3) such as the Gaussian RKHS. We outline the underlying intuition in Section 5.1, and briefly recall the kernel class here.
@@ 5.1 Why the Laplace Kernel? A First Indication
We aim to understand when stationary points of $\mathcal{J}$ recover the relevant features as posed in $(\mathsf{Q4})$. As we will see, the answer depends crucially on the choice of kernel $K$, which fundamentally shapes the optimization landscape.
To build intuition, we consider a simpler diagnostic question:
@@
This question $(\mathsf{Q5})$ serves as a basic test of whether the kernel supports feature detection. Question (i) is the most basic: it asks whether the trivial solution $\beta=0$ can be ruled out a stationary point. Question (ii) is stronger—it asks whether all stationary points capture some relevant feature. The two questions in $(\mathsf{Q5})$ are in fact equivalent under the additional assumptions of Theorem 7, since Theorem 7 ensures that every stationary point satisfies $\mathop{\rm supp}(\beta)\subseteq S_{*}$.
@@ 5.1.1 Comparison to Radial Basis Kernels
$\mathcal{J}(se_{i},\lambda)-\mathcal{J}(0,\lambda)=\int_{0}^{s}\mathrm{D}\mathcal{J}(te_{i},\lambda)[e_{i}]dt=\int_{0}^{s}tG_{i}(t)dt=\frac{s^{2}}{2}G_{i}(0)+o(s^{2})~~(s\to 0^{+}).$
@@ 5.2.3 Error Control
$\begin{split}&\int_{0}^{\infty}\int_{\mathbb{R}^{d}}|\mathbb{E}[R_{\beta,\omega,\lambda}(X,Y)]|^{2}q_{t}(\omega)\,d\omega\,\mu(dt)\\ &=\mathbb{E}\left[r_{\beta,\lambda}(X,Y)r_{\beta,\lambda}(X^{\prime},Y^{\prime})\cdot\int_{0}^{\infty}\int_{\mathbb{R}^{d}}e^{-2\pi i\langle\omega,\beta\circ(X-X^{\prime})\rangle}q_{t}(\omega)\,d\omega\,\mu(dt)\right]\\ &=\mathbb{E}\left[r_{\beta,\lambda}(X,Y)r_{\beta,\lambda}(X^{\prime},Y^{\prime})K(\beta\circ X,\beta\circ X^{\prime})\right].\end{split}$
@@
$\lambda\cdot(\mathrm{D}\mathcal{J}(\beta,\lambda)[e_{i}]-\mathcal{S}_{i}(\beta,\lambda))=\Delta_{1}+\Delta_{2}-\Delta_{3}$
@@
$\begin{split}\Delta_{1}&=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\mathbb{E}[R_{\beta,\omega,\lambda}(X,Y)]\mathbb{E}[\overline{R_{\beta,\omega,\lambda}(X^{\prime},Y^{\prime})}|X_{i}-X_{i}^{\prime}|]~tq_{t}(\omega)d\omega\mu(dt)\\ \Delta_{2}&=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\mathbb{E}[\overline{R_{\beta,\omega,\lambda}(X,Y)}]\mathbb{E}[R_{\beta,\omega,\lambda}(X^{\prime},Y^{\prime})|X_{i}-X_{i}^{\prime}|]~tq_{t}(\omega)d\omega\mu(dt)\\ \Delta_{3}&=\int_{0}^{\infty}\int_{\mathbb{R}^{d}}\mathbb{E}[R_{\beta,\omega,\lambda}(X,Y)]\mathbb{E}[\overline{R_{\beta,\omega,\lambda}(X^{\prime},Y^{\prime})}]\mathbb{E}[|X_{i}-X_{i}^{\prime}|]~tq_{t}(\omega)d\omega\mu(dt)\end{split}.$
The measure $tq_{t}(\omega)d\omega\mu(dt)$ is finite with $\iint tq_{t}(\omega)d\omega\mu(dt)\leq M_{\mu}\iint q_{t}(\omega)d\omega\mu(dt)=M_{\mu}$. Applying the Cauchy Schwarz inequality and invoking Lemmas 18 and 19, we obtain that the error terms $|\Delta_{1}|$ and $|\Delta_{2}|$ are bounded by $C\sqrt{\lambda}$. Applying Lemma 18, we get $|\Delta_{3}|$ is bounded by $C\lambda$. These constants $C<\infty$ depend only on $M_{X},M_{Y},M_{\mu}$.
@@ 5.3.1 Decomposition of Directional Derivative under Functional ANOVA
In particular, the covariance term on the left-hand-side is independent of $\lambda$.
@@ 6 Open Problems
  • (Dependence) In the context of predictive feature recovery, our results assume independence between features. As explained earlier, this independence is mainly used for attributing prediction signals to individual features and feature subsets. Extending the analysis to dependent designs is particularly interesting.
@@ A.2.2 The functional $\mathfrak{L}$ and its Differentiability
$\mathrm{D}\mathfrak{L}(g,\beta,\lambda)[v]:=\lim_{s\to 0^{+}}\frac{1}{s}(\mathfrak{L}(g,\beta+sv,\lambda)-\mathfrak{L}(g,\beta,\lambda)).$ (A.3)